Use the quadratic formula to determine the x-intercepts (if any) of the following function. Then evaluate the function for several values of x.
f(x) = x2 + 2

Answer by rapaljer(4671)

The graph has NO x-intercepts, since x^2+ 2 cannot equal zero.

f(0)=2
f(-1)= 1+2=3
f(-2)=4+2=6
f(1)=1+2=3
f(2)=4+2=6

Dr. Robert J. Rapalje, Retired